Explain the increase in pressure of a gas when its volume is decreased at a constant temperature.


1
Expert's answer
2021-04-18T19:34:45-0400

According to combined gas law:


"\frac{p_1V_1}{T_1}=\frac{p_2V_2}{T_2},\\\space\\
T_1=T_2,\\\space\\
p_1V_1=p_2V_2,\\\space\\
p_2=p_1\frac{V_1}{V_2}."

The volume is decreased, which means that

"V_2<V_1,\rightarrow\frac{V_1}{V_2}>1,"

so, the pressure increases.